A body hanging from a spring is set into motion and the period of oscillation is to be 0.8sec. After the body has come to rest it is removed. How much shorter will the restoring force be when it comes to rest?
1
Expert's answer
2019-11-08T07:27:23-0500

Time period is given by-


"T=2\pi\sqrt{\frac{m}{k}}"

T=0.8 sec

Then,


"0.64=4\pi^2\times\frac{m}{k}\\or\\k=61.68m"

Restoring force =

"kx=61.68mx\\where,\\m=mass\\x= maximum \ displacement \ of\ body"